News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 13 December 1999 9.00am EST

Summary: SI Safing patch uplinked and tested successfully, operations continued nominally.

Over the weekend, Chandra passed through the radiation belts and completed observations of M31 to search for Transients and A2261 with ACIS-I, A2199 and A1835 with ACIS-S, a long observation of NGC 5548 with the HRC-S and LETG, and an observation of Cas A with HRC-S. (Please note that the NGC 1068 observation reported in the 12/10 and 12/8 status reports was observed as planned with ACIS-S and not HRC-S as reported.)

Following radiation belt exit and ACIS CTI measurement at 2:45pm 12/13, the SI Safing patch uplink and test procedure was executed. The software patches were uplinked to the spacecraft during a real-time procedure starting at 2.45pm and completing at 7.30pm. The patches modify the spacecraft safing actions to place the ACIS and HRC in protected configurations in the event of a high radiation environment, unexpected halt of the daily load or safing action such as bright star hold, normal sun mode or safe mode. The test procedure was run to check out the transition to bright star hold, to normal sun mode and to verify correct action in the event of high radiation triggers from the radiation monitor (EPHIN). The test was completed successfully and the modified safing actions are now in place on the A-side computer. Following final assessment of the dump telemetry the B-side patches will be uploaded. Two ground system problems were reported during the test.

During the next 48 hours Chandra will observe TON-S-180 with ACIS-S/LETG, and LP 944-20 with ACIS-S and ACIS-S/LETG.

The spacecraft systems continued to operate nominally.





Return to Status Reports